answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What does major events mean?

Major events means a really important event that happened or is going to happen

How many pips in a deck of cards?

348

the number of pips on cards 2-10= number on the card plus two on opposite corners

(4+5+6+7+8+9+10+11+12)= 72

J, Q, and K each have 4 (2+ 2 in the corners)= 12

A has 3 (1+ 2 in the corners)= 3

pips in each suit= (72+12+3)= 87

multiply by four suits

total pips= 87*4= 348

Is it possible to die from boredom?

When we got not any time for deciding what to do, we only finally die if we cannot find the safest thing to do.

We very probably do not do it when this happens if w a very continuous thing and never find any ways for changing such things when they happen in a very wrong way

How many multiples of 3 are in a deck of cards?

I dont know Good answer right?

----------------------------------------------------

But I do know! The cards which have multiples of 3 are 3, 6. and 9, so there are 3 in each suit but 12 in the whole pack.

How do you use probable cause in a sentence?

Lightening was the probable cause of the power shutdown.

It is probable that the economic situation will deteriorate.

I think lightening is the cause of the bad power .

What is Keynes's theory on probability?

Best Answer

Copy

This answer is very simple .Keynes's theory is a non additive,non numerical , .interval valued approach.It was extended by Keynes to cover decision weights in chapter 26.It also covers additive,numerical, probability if the weight,w,defined on the unit interval [0,1],equals 1.It can easily deal with ordinal probability by the use of standard conditional probability.

What are the names of Jimmie Vaughan's twins?

The twins names are Jenna and Eva, after their two Grandmothers. The twins names are Jenna and Eva, after their two Grandmothers. The twins names are Jenna and Eva, after their two Grandmothers. The twins names are Jenna and Eva, after their two Grandmothers.

What are fraternal twins?

Fraternal twins are two siblings that have the same birth day, but do not look identical.

Basketball player Chauncey Billups of the Detroit Pistons makes free throw shots 88 percent of the time Find the probability that he misses his first shot and makes the second?

10.56%

The probability that he misses his first shot is 12%. The probability that he makes the second shot is 88%. The probability of missing the first shot and making the second shot is 12% * 88%, or 0.12*0.88*100.

The first letter of each month is written on a card and the cards are placed in a bag. What isthe probability of pulling out a J, placing it back in the bag, and then drawing an M?

So, to start with, the probability of drawing a J is 3/12 since there is January, July, and June. So there's a 25% chance of getting J first.

There are 2 months starting with M, March and May, so by itself there is a 2/12 chance of that happening, or a roughly 17% chance.

However, since we only care about drawing an M after drawing a J, we have a 3/12 chance of the first draw, and the other 9/12 possibilities we don't care about because they've already failed. So out of the 3/12 times where we draw a J, we then have a 2/12 chance of drawing an M.

So 3/12 * 2/12 = 6/144, about a 4% chance of drawing these letters in this order.

Specifically, 0.04166666

What is the probability that you do not pick a picture card?

There are 12 picture cards and 38 non-picture cards in a deck of 52 cards. The probability that you do not pick a picture card is therefore 38/52 = 19/26.

What is the probability that a randomly chosen month has 30 days?

there are 4 monthes that has 30 days each year

there are 12 monthes each year.

so the probabillity for picking a month that has 30 days

in it is 4/12

How to find Median of probability distribution?

how do i find the median of a continuous probability distribution

Function of a basophil?

Basophils appear in many specific kinds of inflammatory reactions, particularly those that cause allergic symptoms. Basophils contain anticoagulant heparin, which prevents blood from clotting too quickly. They also contain the vasodilator histamine, which promotes blood flow to tissues.

How fast does jennys toy car roll if it can roll 16 inches in 2 seconds?

16 inches/2 sec = 8 inches per second

thats is 8/12 = 2/3 feet per second and since 88 feet per sec = 60 mph that is

2/3/88 x 60 = less than 1/2 mile per hour

Trending Questions
The first organisms were most probably? Number that describes a certain event will occur? The process of determining whether information has already been orignially classified is? A large bag of marbles contains 57 blue marbles and 74 red marbles and 69 yellow marbles What percent of the marbles are red? If you draw a card from a normal deck of cards what is the probability that you will draw an even number or red card? How many reds in a deck of 52 cards? What is the mean mode median standard deviation and standard error of 10 24 35 44 10 and 35? In a bag there 2 red 3yellow 4green and 6 blue marbles What is the probability of P Blue? How many cards are in a standard deck of playing? When rolling a cube what is the probability of rolling a 1 and a 6 at the same time? These bag contains 6 white ball 3 red balls and one blue ballif one ball is drawn from the bag what is the probability it will be white? What does it mean when an event has a probability of 1? What is the definition for Probability? What is the z-score of z0.07? What is the sample space of rolling a die? How do you calculate the probability for a group of several independent events? What is the probability of spinning a number greater than 5 on a spinner numbered 1 to 8 and tossing a tail on a coin? What is the probability of not getting an even number on the spinner? What is Critical cross field outcomes? How many outcomes are there for flipping a coin and spinning a spinner that has 6 different colored sections?